# Lipschitz-free spaces and Bossard’s reduction argument

### v1

@@ Abstract.
Motivated by recent work of E. Basset, G. Lancien and A. Procházka, we use the reduction argument of Bossard to prove two results: if a separable Banach space is isomorphically universal for the class of Lipschitz-free spaces over the countable complete discrete metric spaces then it is isomorphically universal for the class of separable Banach spaces, and if a complete separable metric space is Lipschitz universal for the same class of metric spaces then it is Lipschitz universal for all separable metric spaces.
@@
Lipschitz-free space, universality, descriptive set theory
@@ 1. Introduction
Let $M$ be a metric space with a base point denoted by 0. Let $\operatorname{Lip}_{0}(M)$ denote the Banach space of real-valued Lipschitz functions that vanish at 0, endowed with the norm that assigns to each $f\in\operatorname{Lip}_{0}(M)$ its optimal Lipschitz constant. Recall that the map ${\delta}:{M}\to{\operatorname{Lip}_{0}(M)^{*}}$ given by $\langle{f},{\delta(x)}\rangle=f(x)$, $x\in M$, $f\in\operatorname{Lip}_{0}(M)$, is an isometric embedding, and that we can define the Lipschitz-free space (hereafter free space) $\mathcal{F}({M})$ over $M$ as the closed linear span of the image $\delta(M)\subset\operatorname{Lip}_{0}(M)^{*}$ (where $\delta(M\setminus\{0\})$ is linearly independent).
@@
The first example of an uncountable family of separable metric spaces whose free spaces are pairwise non-isomorphic was given in [hajek:lancien:pernecka:16]. These metric spaces are all homeomorphic to the Cantor set $2^{\omega}$; the fact that their corresponding free spaces are pairwise non-isomorphic relies on the theorem of Johnson and Szankowski which states that there is no separable Banach space complementably universal for all separable Banach spaces.
Two more examples of such uncountable families are given in [basset:24, basset:lancien:prochazka:25]. In [basset:24], each metric space in the family is countable and complete, while in [basset:lancien:prochazka:25] each one is countable complete and discrete. The free spaces corresponding to the first family are non-isomorphic because the weak-fragmentability (or weak Szlenk) index of each one is distinct (and countable). The same holds in the case of the second family, this time with respect to the dentability index (not to be confused with the $w^{*}$-dentability index).
@@
I heard about the results in [basset:lancien:prochazka:25] in a talk given by E. Basset at the Banff International Research Station workshop on Functional and Metric Analysis and their Interactions, held at the Institute of Mathematics, University of Granada, Spain, in May 2025. It occurred to me that it might be possible to use descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p. 31), to build on their universality results. I would like to thank E. Basset, G. Lancien and A. Procházka for the fascinating exchange that ensued. I would also like to thank G. Godefroy for very interesting comments about the wider area. The two main results of the paper are given below. The second one was prompted by a question asked by G. Lancien.
@@
We prove some refinements of these results and make some related minor observations along the way. The application of descriptive set theory to Banach space theory can be traced to some seminal papers of Bourgain, e.g. [bourgain:80]. It was formalised later by Bossard in another seminal paper [bossard:02] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g. [dodos:10] and references therein, and [cuth:doucha:kurka:22b] and related papers for more recent developments. We speculate that techniques from descriptive set theory could be used to prove more results about free spaces in the future. Finally, I should stress that I have left out most of the results in [basset:24, basset:lancien:prochazka:25] from this introduction purely in the interests of brevity, and I recommend that the interested reader should consult these papers in full.
@@ 2. Notation and preliminaries
We use standard Banach space notation throughout. The closed unit ball of a Banach space $X$ is denoted $B_{X}$. Given Banach spaces $X$ and $Y$, we write $X\hookrightarrow Y$ and $X\equiv Y$ when $X$ is linearly isomorphic to a subspace of $Y$ and $X$ is linearly isometric to $Y$, respectively. We also use these symbols in the context of metric spaces $M$ and $N$, with “Lipschitz” substituted for “linear”. By [godefroy:kalton:03]*Lemmas 2.2 and 2.3 (or [weaver:18]*Theorems 3.6 and 3.7), if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense). We say that the Banach space $X$ is isomorphically universal for a given class $\mathcal{C}$ of Banach spaces if $Y\hookrightarrow X$ whenever $Y\in\mathcal{C}$; Lipschitz universality of metric spaces is defined analogously.
For concepts pertaining to descriptive set theory, we follow [kechris:95]. The sets of natural numbers, countable ordinals and infinite sequences of natural numbers are denoted $\omega$, $\omega_{1}$ and $\omega^{\omega}$, respectively. We identify the power set of a set $\Gamma$ with $2^{\Gamma}$. We denote by $\mathbf{\Sigma}^{0}_{\alpha}$ and $\mathbf{\Pi}^{0}_{\alpha}$ the additive and multiplicative Borel classes of order $\alpha<\omega_{1}$ in a Polish space, so that $F_{\sigma}$, $G_{\delta}$ and $F_{\sigma\delta}$-sets are $\mathbf{\Sigma}^{0}_{2}$, $\mathbf{\Pi}^{0}_{2}$ and $\mathbf{\Pi}^{0}_{3}$-sets, and so on. Analytic and coanalytic subsets of Polish spaces are called $\mathbf{\Sigma}^{1}_{1}$ and $\mathbf{\Pi}^{1}_{1}$-sets, respectively. Given subsets $A\subset X$, $B\subset Y$ of Polish spaces $X$ and $Y$, we say that $A$ is reducible to $B$ if there is a continuous function ${f}:{X}\to{Y}$ such that $A=f^{-1}(B)$. If $\mathbf{\Gamma}(X)$ is a class of subsets of a Polish space $X$, a subset $B\subset Y$ of a Polish space $Y$ is called $\mathbf{\Gamma}$-complete if $B\in\mathbf{\Gamma}(Y)$ and $A$ is reducible to $B$ whenever $A\in\mathbf{\Gamma}(X)$ and $X$ is Polish and zero-dimensional.
@@
Next, we set up an appropriate descriptive set-theoretic framework for the class of separable infinite-dimensional free spaces. We shall consider the Polish space $0,\infty)^{\omega\times\omega}$ endowed with the pointwise topology, and the set $\mathcal{R}\subset[0,\infty)^{\omega\times\omega}$ of metrics on $\omega$. This space was introduced by Vershik in [vershik:98] and features again in e.g. [vershik:04] and [clemens:12] (in these works $\mathcal{R}$ is defined instead as the set of all pseudo-metrics on $\omega$, that is, functions $d\in[0,\infty)^{\omega\times\omega}$ that are metrics except that it is possible to have distinct $i,j\in\omega$ satisfying $d(i,j)=0$). It is easily seen that $\mathcal{R}$ is a $G_{\delta}$ in $0,\infty)^{\omega\times\omega}$ and is thus Polish as well. This space can be used to code the class of complete separable infinite metric spaces. Given such a space $(M,\rho)$, fix a dense sequence of distinct points $(x_{i})\subset M$ and consider $d\in\mathcal{R}$ given by $d(i,j)=\rho(x_{i},x_{j})$, $i,j\in\omega$; then $(M,\rho)\equiv M_{d}$, where $M_{d}$ is the completion of $(\omega,d)$. Conversely any such completion is separable and infinite. Because a free space is separable and infinite-dimensional if and only if its underlying metric space is separable and infinite, and because the free space over a metric space is isometric to that of its completion, $\mathcal{R}$ can be seen as a natural coding of the class of separable infinite-dimensional free spaces. As a change of base point does not alter the isometric structure of the corresponding free space, we shall always regard $0\in\omega$ as the base point.
@@
Bossard coded the class of separable infinite-dimensional Banach spaces as a Borel subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure. The authors of [cuth:doucha:kurka:22a, cuth:doucha:kurka:22b] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$. First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm. Endowed with the pointwise topology, $\mathcal{B}$ is a $G_{\delta}$ in $0,\infty)^{V}$ and is thus Polish. Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$. Given a dense linearly independent sequence $(x_{i})\subset X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$. Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space. We will work with $\mathcal{B}$ instead of $\mathcal{SB}$ because, in our opinion, it is easier to use and more compatible with $\mathcal{R}$. While not directly relevant to this paper, another advantage of using $\mathcal{B}$ is that Baire category arguments can be applied, e.g. the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [cuth:doucha:kurka:22b]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space. According to [cuth:doucha:kurka:22a]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$. It follows easily from this and [bossard:02]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
@@
Concerning Polish versus Effros-Borel structures, we should mention that it is possible to code the class of complete infinite separable metric spaces using the set of closed infinite subsets of the Urysohn universal separable metric space $\mathbb{U}$, again equipped with the Effros-Borel structure, see e.g. [gao:09, gao:kechris:03]. However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above. Analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [vershik:04]*Theorems 1 and 3. Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [fonf:wojtaszczyk:08], could be regarded as the generic separable free space. A immediate consequence of this and [fonf:wojtaszczyk:08]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual. We mention this because Polish spaces of metrics have been used elsewhere to study approximation properties in free spaces [talimdjioski:23, smith:talimdjioski:24, talimdjioski:25].
3. The main results

### v2

@@ Abstract.
We set up a descriptive set-theoretic framework to study Lipschitz-free spaces and use the reduction argument of Bossard to prove several results. We prove two universality results: if a separable Banach space is isomorphically universal for the class of Lipschitz-free spaces over the countable complete discrete metric spaces then it is isomorphically universal for the class of separable Banach spaces, and if a complete separable metric space is Lipschitz universal for the same class of metric spaces then it is Lipschitz universal for all separable metric spaces. We also show that there exist countable complete discrete metric spaces whose Lipschitz-free spaces fail the bounded approximation property and are thus not isomorphic to any dual Banach space. Finally, we calculate the descriptive complexity of the classes of separable Banach spaces and separable Lipschitz-free spaces having the approximation property.
@@
Lipschitz-free space, universality, descriptive set theory, approximation property, bounded approximation property
@@ 1. Introduction
Let $M$ be a metric space with a base point denoted by 0. Let $\operatorname{Lip}_{0}(M)$ denote the Banach space of real-valued Lipschitz functions that vanish at 0, endowed with the norm that assigns to each $f\in\operatorname{Lip}_{0}(M)$ its optimal Lipschitz constant. Recall that the map ${\delta}:{M}\to{\operatorname{Lip}_{0}(M)^{*}}$ given by $\langle{f},{\delta(x)}\rangle=f(x)$, $x\in M$, $f\in\operatorname{Lip}_{0}(M)$, is an isometric embedding, and that we can define the Lipschitz-free space (hereafter free space) $\mathcal{F}({M})$ over $M$ as the closed linear span of the image $\delta(M)\subseteq\operatorname{Lip}_{0}(M)^{*}$ (where $\delta(M\setminus\{0\})$ is linearly independent).
@@
The first example of an uncountable family of separable metric spaces whose free spaces are pairwise non-isomorphic was given in [hajek:lancien:pernecka:16]. These metric spaces are all homeomorphic to the Cantor set $2^{\omega}$. Two more uncountable families of this kind are given in [basset:24, basset:lancien:prochazka:25]. In [basset:24], each metric space in the family is countable and complete, while in [basset:lancien:prochazka:25] each one is countable complete and discrete. The free spaces corresponding to the first family are non-isomorphic because the weak-fragmentability (or weak Szlenk) index of each one is distinct (and countable). The same holds in the case of the second family, this time with respect to the dentability index (not to be confused with the $w^{*}$-dentability index).
@@
In this paper we use a tool from descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p. 31), to build on this universality result.
@@
Another enduring strand of this research has been to identify which free spaces have or fail one of Grothendieck’s classical approximation properties. Recall that a Banach space $X$ is said to have the approximation property (AP) if, given compact $K\subseteq X$ and $\varepsilon>0$, there exists a bounded linear finite-rank operator ${T}:{X}\to{X}$ such that $\left\|{Tx-x}\right\|\leqslant\varepsilon$ whenever $x\in K$. If there exists $\lambda\geqslant 1$ such that the operator $T$ can always be chosen to satisfy $\left\|{T}\right\|\leqslant\lambda$, then $X$ is said to have the $\lambda$-bounded approximation property ($\lambda$-BAP). We say that $X$ has the bounded approximation property (BAP) if it has the $\lambda$-BAP for some $\lambda$, and the metric approximation property (MAP) if it has the $1$-BAP.
The literature on free spaces and approximation properties is too abundant to be summarised in this short paper; we invite the interested reader to consult [godefroy:24] and the introduction of [smith:talimdjioski:23] and references therein for more information. We only need to recall two results for our purposes. First, it follows immediately from [godefroy:kalton:03]*Theorem 3.1 that if a separable Banach space $X$ fails the AP then so does $\mathcal{F}({X})$. Second, [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8 implies that $\mathcal{F}({M})$ has the AP whenever $M$ is countable discrete and complete. To the best of my knowledge, to date there has been no example of a complete separable p1u metric space whose free space has the AP but fails the BAP. The next result shows that there are such examples and, moreover, they can be chosen to be countable complete and discrete. Again, this result is based on Bossard’s argument. Another crucial ingredient is the fact that, relative to the established descriptive set-theoretic frameworks for separable Banach spaces (see Section 2), the class of separable Banach spaces that satisfies the BAP is Borel (though being analytic would have sufficed) [ghawadrah:17].
Theorem 1.3.
There exist countable complete discrete metric spaces $M$ such that $\mathcal{F}({M})$ fails the BAP. Consequently, $\mathcal{F}({M})$ is not isomorphic to a dual space.
This result can be contrasted with [aliaga:gartand:petitjean:prochazka:22]*Theorem 3.2 which states, among other things, that $\mathcal{F}({M})$ is isometric to a dual space whenever $M$ is a p1u metric space that is proper, i.e. its closed balls are compact. It can also be regarded in the light of Kalton’s problem of whether $\mathcal{F}({M})$ has the BAP whenever $M$ is uniformly discrete; see [kalton:04]*p. 185 and e.g. [godefroy:24]*Problem 4.1 for more context. In this vein, there exists a uniformly discrete space whose free space is not isometric to a dual space (though this particular example has the MAP) [garcia-lirola:petitjean:prochazka:ruedazoca:18]*Example 5.8.
We prove the results above using Lemma 3.10, which is one of our two main tools. The other main tool, Lemma 3.9, will be needed for our final main result, Theorem 4.1, which shows that the classes of separable Banach spaces and separable free spaces having the AP are, relative to the appropriate descriptive set-theoretic interpretations, complete coanalytic sets. (We remark that this is not the first time Polish spaces of metrics have been used to study free spaces and approximation properties; see [talimdjioski:23, smith:talimdjioski:24, talimdjioski:25].)
@@
We prove these theorems in Sections 3 and 4, and make some related observations along the way. The application of descriptive set theory to Banach space theory can be traced to some seminal papers of Bourgain, e.g. [bourgain:80]. It was formalised later by Bossard in another seminal paper [bossard:02] and has since developed into a rich and beautiful subfield that boasts many powerful techniques and results, see e.g. [dodos:10] and references therein, and [cuth:doucha:kurka:22b] and related papers for more recent developments. I speculate that techniques from descriptive set theory could be used to prove more results about free spaces in the future.
I heard about the results in [basset:lancien:prochazka:25] in a talk given by E. Basset at the Banff International Research Station workshop on Functional and Metric Analysis and their Interactions, held at the Institute of Mathematics, University of Granada, Spain, in May 2025. It occurred to me that it might be possible to use descriptive set theory, specifically Bossard’s reduction argument (see [dodos:10]*p. 31), to build on their universality results. I would like to thank E. Basset, G. Lancien and A. Procházka for the fascinating exchanges that ensued. In particular, Theorem 1.2 was prompted by a question asked by G. Lancien. I would also like to thank M. Cúth, G. Godefroy and A. Rueda Zoca for several remarks which helped me when I was preparing this paper. Finally, I should stress that I have left out most of the results in [basset:24, basset:lancien:prochazka:25] from this introduction purely in the interests of brevity, and I recommend that the interested reader should consult these papers in full.
@@ 2. Notation and preliminaries
We use standard Banach space notation throughout. The closed unit ball of a Banach space $X$ is denoted $B_{X}$. Given Banach spaces $X$ and $Y$, we write $X\hookrightarrow Y$ and $X\equiv Y$ when $X$ is linearly isomorphic to a subspace of $Y$ and $X$ is linearly isometric to $Y$, respectively. We also use these symbols in the context of metric spaces $M$ and $N$, with “Lipschitz” substituted for “linear”. By [weaver:18]*Theorems 3.6 and 3.7, if $M\hookrightarrow N$ (in the Lipschitz sense) then $\mathcal{F}({M})\hookrightarrow\mathcal{F}({N})$ (in the linear sense). We say that the Banach space $X$ is isomorphically universal for a given class $\mathcal{C}$ of Banach spaces if $Y\hookrightarrow X$ whenever $Y\in\mathcal{C}$; Lipschitz universality of metric spaces is defined analogously.
For concepts pertaining to descriptive set theory, we follow [kechris:95]. The sets of natural numbers, countable ordinals and infinite sequences of natural numbers are denoted $\omega$, $\omega_{1}$ and $\omega^{\omega}$, respectively. We identify the power set of a set $\Gamma$ with $2^{\Gamma}$. We denote by $\mathbf{\Sigma}^{0}_{\alpha}$ and $\mathbf{\Pi}^{0}_{\alpha}$ the additive and multiplicative Borel classes of order $\alpha<\omega_{1}$ in a Polish space, so that $F_{\sigma}$, $G_{\delta}$ and $F_{\sigma\delta}$-sets are $\mathbf{\Sigma}^{0}_{2}$, $\mathbf{\Pi}^{0}_{2}$ and $\mathbf{\Pi}^{0}_{3}$-sets, and so on. Analytic, coanalytic and Borel subsets of Polish spaces are called $\mathbf{\Sigma}^{1}_{1}$, $\mathbf{\Pi}^{1}_{1}$ and $\mathbf{\Delta}^{1}_{1}$-sets, respectively. Given subsets $A\subseteq X$, $B\subseteq Y$ of Polish spaces $X$ and $Y$, we say that $A$ is reducible to $B$ if there is a continuous function ${f}:{X}\to{Y}$ such that $A=f^{-1}(B)$. If $\mathbf{\Gamma}(X)$ is a class of subsets of a Polish space $X$, a subset $B\subseteq Y$ of a Polish space $Y$ is called $\mathbf{\Gamma}$-hard if $A$ is reducible to $B$ whenever $A\in\mathbf{\Gamma}(X)$ and $X$ is Polish and zero-dimensional, and $\mathbf{\Gamma}$-complete if $B$ is $\mathbf{\Gamma}$-hard and in addition belongs to $\mathbf{\Gamma}(Y)$.
@@
Next, we set up an appropriate descriptive set-theoretic framework for the class of separable infinite-dimensional free spaces. We shall consider the Polish space $0,\infty)^{\omega\times\omega}$ endowed with the pointwise topology, and the set $\mathcal{R}\subseteq[0,\infty)^{\omega\times\omega}$ of metrics on $\omega$. This space was introduced by Vershik in [vershik:98] and features again in e.g. [vershik:04] and [clemens:12] (in these works $\mathcal{R}$ is defined instead as the set of all pseudo-metrics on $\omega$, that is, functions $d\in[0,\infty)^{\omega\times\omega}$ that are metrics except that it is possible to have distinct $i,j\in\omega$ satisfying $d(i,j)=0$). It is easily seen that $\mathcal{R}$ is a $G_{\delta}$ in $0,\infty)^{\omega\times\omega}$ and is thus Polish as well. This space can be used to code the class of complete separable infinite metric spaces. Given such a space $(M,\rho)$, fix a dense sequence of distinct points $(x_{i})\subseteq M$ and consider $d\in\mathcal{R}$ given by $d(i,j)=\rho(x_{i},x_{j})$, $i,j\in\omega$; then $(M,\rho)\equiv M_{d}$, where $M_{d}$ is the completion of $(\omega,d)$. Conversely any such completion is separable and infinite. Because a free space is separable and infinite-dimensional if and only if its underlying metric space is separable and infinite, and because the free space over a metric space is isometric to that of its completion, $\mathcal{R}$ can be seen as a natural coding of the class of separable infinite-dimensional free spaces. As a change of base point does not alter the isometric structure of the corresponding free space, we shall always regard $0\in\omega$ as the base point.
@@
Bossard coded the class of separable infinite-dimensional Banach spaces as a $\mathbf{\Delta}^{1}_{1}$ subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure. Following a programme initiated in [godefroy:saint-raymond:18], the authors of [cuth:doucha:kurka:22a, cuth:doucha:kurka:22b] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$. First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm. Endowed with the pointwise topology, $\mathcal{B}$ is a $G_{\delta}$ in $0,\infty)^{V}$ and is thus Polish. Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$. Given a dense linearly independent sequence $(x_{i})\subseteq X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$. Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space. We will work with $\mathcal{B}$ instead of $\mathcal{SB}$ because, in our opinion, it is easier to use and more compatible with $\mathcal{R}$. While not directly relevant to this paper, another advantage of using $\mathcal{B}$ is that Baire category arguments can be applied, e.g. the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [cuth:doucha:kurka:22b]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space. According to [cuth:doucha:kurka:22a]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$. It follows easily from this and [bossard:02]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
@@
Concerning Polish versus Effros-Borel structures, we should mention that it is possible to code the class of complete infinite separable metric spaces using the set of closed infinite subsets of the Urysohn universal separable metric space $\mathbb{U}$, again equipped with the Effros-Borel structure, see e.g. [gao:09, gao:kechris:03]. However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above. Also, analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [vershik:04]*Theorems 1 and 3. Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [fonf:wojtaszczyk:08], could be regarded as the generic separable free space. A immediate consequence of this and [fonf:wojtaszczyk:08]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual. This observation complements Theorem 4.1.
3. Three main results